# Graph Masked Autoencoder for Spatio-Temporal Graph Learning

### v1

@@ IV-A3 Implementation Details
Table III: Overall performance comparison in crime prediction on both Chicago and NYC datasets.
Chicago New York City
Model Theft Battery Assault Damage Burglary Larceny Robbery Assault
MAE MPAE MAE MPAE MAE MPAE MAE MPAE MAE MPAE MAE MPAE MAE MPAE MAE MPAE
Node2vec 1.1378 0.9862 1.7655 0.8970 1.9631 0.9714 1.9015 0.9657 4.9447 0.8092 0.7272 0.6532 1.0566 0.8040 1.2411 0.9967
GCN 1.1065 0.9643 1.3012 0.8094 1.5431 0.8094 1.5031 0.8056 4.6993 0.7912 0.49994 0.4178 1.0655 0.8004 1.2407 0.9890
GAT 1.1123 0.9759 1.3215 0.8344 1.5892 0.8241 1.5387 0.8277 4.7055 0.7944 0.5023 0.4019 1.0653 0.8027 1.2403 0.9949
GraphSage 1.1231 0.9790 1.3574 0.8561 1.6016 0.8563 1.5761 0.8432 4.7313 0.8066 0.5213 0.4314 1.0719 0.8110 1.2418 0.9965
GAE 1.1043 0.9614 1.3065 0.7984 1.5379 0.7914 1.4986 0.8033 4.7013 0.7910 0.5012 0.4289 1.0679 0.8012 1.2405 0.9958
GraphCL 0.9994 0.8992 1.0578 0.6174 1.2938 0.5683 1.2457 0.6181 4.5047 0.7288 0.3834 0.3091 1.0498 0.5961 0.7724 0.6929
RGCL 0.9377 0.8779 1.0356 0.5865 1.2764 0.5457 1.2336 0.5975 4.4712 0.7064 0.3652 0.2832 0.9987 0.5769 0.7512 0.6743
POI 0.9733 0.9341 1.1065 0.7513 1.4089 0.7541 1.4076 0.7697 4.6939 0.7825 0.4969 0.4172 1.0660 0.7970 1.2400 0.9943
HDGE 0.9545 0.9012 1.0887 0.7389 1.3970 0.7217 1.3768 0.7349 4.5658 0.7160 0.4734 0.3930 1.0507 0.6731 1.1551 0.9870
ZE-Mob 1.0983 0.9547 1.3142 0.8236 1.5345 0.8163 1.5138 0.8273 4.7570 0.8013 0.5186 0.4307 1.0722 0.8093 1.1403 0.9975
MV-PN 0.9613 0.9146 1.0946 0.7452 1.4013 0.7393 1.3582 0.7218 4.6329 0.7502 0.4213 0.3708 1.0642 0.7840 1.1091 0.9982
CGAL 0.9589 0.9014 1.0897 0.7403 1.3995 0.7345 1.3698 0.7296 4.6013 0.7203 0.4113 0.3651 1.0714 0.7765 1.1009 0.9894
MVURE 0.9365 0.8910 1.0631 0.6957 1.3709 0.6375 1.3037 0.6567 4.5907 0.7144 0.4077 0.3262 1.0578 0.5889 0.8410 0.6943
MGFN 0.9231 0.9015 1.0804 0.5824 1.3016 0.6072 1.2563 0.6503 4.5646 0.7994 0.4285 0.3084 1.0475 0.6310 0.8319 0.7096
STGMAE 0.7860 0.7365 1.0377 0.4207 1.1830 0.4214 0.9252 0.4596 4.3095 0.6404 0.2958 0.1924 0.8788 0.3789 0.6654 0.5537
IV-A4 Baselines
We use 15 baselines from three research lines namely, graph representation approaches, graph contrastive learning models, and spatial-temporal region representation methods for comparison in order to thoroughly assess our STGMAE model.
@@ IV-A4 Baselines
  • Node2vec [34] encodes graph structural data using a random walk-based skip-gram.
  • GCN [28] carries out the convolution-based message transfer along the edges between neighboring nodes for refinement.
  • GraphSage [35] is a graph neural architecture that permits the accumulation of information from the sub-graph structures.
  • GAE [17] maps node into latent embedding space by the Graph Auto-encoder using the input reconstruction objective.
  • GAT [36] varies the degrees of significance among nearby nodes, and improves GNNs’ capacity to discriminate between diverse inputs. To optimize the performance of GNN-based approaches, the number of information propagation layers is adjusted from a range of {1,2,3,4}.
@@
  • GraphCL [37] is based on the maximizing of mutual knowledge and generates many contrastive viewpoints for augmentation. The goal is to ensure embedding consistency across various connected views via the InfoNCE loss.
  • RGCL [38] does data augmentation based on the created rationale generator, is used to do so. Different GCL models’ temperature parameters have been fine-tuned to manage the contrastive augmentation in order to allow for fair comparison.
@@
  • POI [29] uses TF-IDF tokenization with the provided POI matrix and the POI properties to represent spatial regions.
  • HDGE [22] creates a crowd flow graph using human trajectory data and embeds regions into latent vectors to protect the structural data of the network and capture mobility patterns.
  • ZE-Mob [23] uses region correlations to create embeddings while taking into account human movement and taxi moving traces based on multi-view data sources.
  • MV-PN [26] models intra-regional and inter-regional correlations with an encoder-decoder network.
  • CGAL [25] is a technique for graph-regularized adversarial learning that uses pairwise graph relationships to embed regions in latent space via the graph attention neural network.
  • MVURE [24] makes use of the graph attention mechanism to simulate region correlations with the built-in properties of the region and data on human mobility.
  • AutoST [15] is a method that utilizes contrastive self-supervised learning method via multi-view graph to model region representation learning problem on traffic and crime prediction tasks.
  • MGFN [27] aggregates information for both intra-pattern and inter-pattern patterns, and encodes region embeddings with multi-level cross-attention without region function information.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Results on NYC and CHI crime for four crime types w.r.t different data density degrees.
IV-B Effectiveness Evaluation (RQ1)
Table II shows a performance comparison of all methods on three tasks (crime prediction, traffic prediction, and house price prediction) in terms of MAE, MAPE, and RMSE.
Crime prediction. We show crime prediction results in Table II and type-specific results in Table III. From Table II, we summarize observations as follows:
  • On both datasets, our framework STGMAE performs best at predicting crimes, indicating the viability of our generative spatio-temporal graph self-supervised learning approach. In specific, we attribute the benefits to: i) STGMAE can capture both the mutli-view area relationships from multi-typed data sources (i.e., POI semantics, human motion traces, geographical coordinates) thanks to masking different types of relations. ii) The adaptive data augmentation way of STGMAE via the mask mechanism on nodes and relations improves the representation ability against noise disturbance and skewed data distribution.
  • While existing approaches for region representation (such as CGAL, MVURE, and MGFN) aim to capture the mobility-based region correlations, they are susceptible to noisy edges in their built mobility-based region graphs. The performance of region representation learning will be hampered by the propagation of information propagation between less-relevant regions. Besides, the spatial dependence modeling in such baselines is constrained by the skewed distribution of mobility data. Our STGMAE   which is complementary to the mobility-based region connections, adds region self-discrimination supervision signals to the region embedding paradigm to solve these limitations.
  • The large enhancement of STGMAE over the compared graph embedding techniques (i.e., GCN and GAT) further supports the significance of generative self-supervised learning paradigm. When a message passes over the edges of a task-irrelevant region in the graph neural paradigm, noise effects will be enhanced. The information compiled from all the nearby mobility-based or POI-based graph regions may be misleading when it comes to encoding the actual underlying crime trends.
Traffic prediction. We present traffic prediction results in Table II with the following observations:
  • The significance of the generative self-supervised learning with the mask mechanism is further supported by the significant performance improvement of STGMAE over the compared graph embedding techniques (such as GCN, GAT, and GAE). In the graph neural paradigm, noise effects will be amplified when messages crosses edges of regions that are unrelated to the task.
@@ IV-B Effectiveness Evaluation (RQ1)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Hyperparameter study on traffic prediction
IV-C Ablation Study (RQ2)
@@
To evaluate the effectiveness of each component of our framework STGMAE, we perform an ablation study and present results w.r.t traffic prediction and crime prediction in Figure 2 and Figure 3. Particularly, we conduct an ablation study on “RP GCN” by replacing GCN with MLP, “w/o Structure Mask”, “w/o Node Mask”, and “w/o Edge Type Mask” by removing masking links or masking nodes.
@@
Analysis on Crime prediction. The figure demonstrates the impact of different components on region representation for crime prediction. “RP GCN” replacing GCN with a 2-layer MLP, contributes the most to performance improvement compared to “w/o Structure Mask”, “w/o Edge Type Mask”, and “w/o Node Mask” GCN captures spatial-temporal graph structure semantics, unlike MLP. “w/o Node Mask” has a larger impact than “w/o Structure Mask” as masking nodes directly affects region representations. Masking different link types positively affects region representations, emphasizing the significance of masking links in the heterogeneous graph. “w/o Edge Type Mask” which doesn’t mask edges, shows a notable performance decline, highlighting the effectiveness of edge type masking in our STGMAE model. Overall, integrating GCN, node masking, and edge-type masking enhances the performance of our STGMAE framework.
@@
Analysis on Traffic prediction. Based on Figure 2, GCN has the greatest impact on improving region representations and traffic prediction, especially for longer time horizons. Masking edge types significantly improves traffic flow prediction accuracy. Masking nodes and graph structure also contribute positively by refining region representations and capturing spatial relationships. Incorporating POI information enhances region representations and enables the encoding of different region relations. Overall, the combination of GCN, edge type masking, node masking, and POI integration contributes to the success of our STGMAE method in learning region representations and achieving accurate traffic predictions.
Type-Specific Crime Prediction Results. We present a comprehensive evaluation of various criminal offense types in the cities of Chicago and New York. The evaluation measures the performance of 16 different approaches in terms of MAE and MAPE. The foundational method for the other 16 approaches is ST-SHN. The results, shown in Table III, clearly demonstrate that our STGMAE approach consistently outperforms all other methods across all crime categories in both cities. This highlights the significant advantages that our region’s embedding learning framework brings to the table. We attribute the success of our approach to the effectiveness of the multi-view region graph’s graph encoding, which extracts valuable regional features for region representation. Additionally, the generative contrastive method, utilizing masking nodes and different types of links in the heterogeneous graph, plays a crucial role. This further validates the effectiveness of generative self-supervised learning in spatio-temporal graph-related applications such as traffic flow prediction, crime prediction and house prediction, leveraging the mask mechanism on nodes and structures.
IV-D More Effectiveness Evaluation Results (RQ1)
@@
Results for the data on home prices in Chicago and New York City are shown in Table II and more analysis details are below. A Lasso regression approach uses the input embeddings from the learned region representations created using various methods.
@@
  • STGMAE consistently performs at the highest level. On the other hand, MVURE and MGFN perform better when using data from the housing markets of New York and Chicago, respectively. This is because they mine human patterns, which connect comparable or nearby places. We also notice that the POI technique performs well because it creates connections between POIs of regions and core shopping regions, which include POIs that are distinct from those in remote towns.
@@
  • Table IV shows results of improvement of existing widely used traffic prediction methods (e.g., GMAN, STGCN) based on several region representation baselines and our method STGMAE. The superior performance further validates the effectiveness of generative self-supervised learning with a masking mechanism for denoising data. In contrast, methods like MVURE and MGFN require significantly more data to uncover human patterns.
Table IV: Overall Performance Comparison in Traffic Prediction Improvement
@@
w/STGMAE 1.8403 2.7032 w/STGMAE 1.2415 3.6375
IV-E Model Robustness Study (RQ3)
@@
We also conduct experiments to test how well our framework STGMAE stands up to data sparsity. To accomplish this, we assess the forecast accuracy of regions with various densities independently. Here, the ratio of non-zero elements (crime occurs) in the region-specific crime occurrence sequence $\mathbf{Z}_{r}$ is used to estimate the density degree of each region. We divide sparse areas with a crime density degree of less than 0.5 into two groups, (0.0, 0.25] and (0.25, 0.5] specifically. Figure 4 presents the evaluation findings. We see that our STGMAE constantly performs better than alternative approaches. Thus, this experiment once more indicates how masking mechanism on nodes and the structure for self-supervised learning into the spatial-temporal region representation learning provides very accurate and reliable crime prediction in all scenarios with varying crime density degrees. To mine human behaviors or movement patterns, existing region representation algorithms (such MVURE and MGFN) require a huge quantity of data, which results in lower performance on the sparse data. Another finding acknowledges that typical graph learning techniques now in use (such as GNN-based approaches including GCN, GAE and GAT) are unable to efficiently build high-quality representations of regions with sparse crime data via the spatial-temporal graph.
@@
The sencond line of Figure 4 present more details about the evaluation findings of the model robustness. We see that STGMAE constantly performs better than alternative approaches. Thus, this experiment once more indicates how masking mechanism on nodes and the structure for self-supervised learning into the spatial-temporal region representation learning provides very accurate and reliable crime prediction in all scenarios with varying crime density degrees. To mine human behaviors or movement patterns, existing region representation algorithms (MVURE and MGFN) require a huge quantity of data, which results in lower performance on the sparse data. Another finding acknowledges that typical graph learning techniques now in use (i.e., GNN-based approaches including GCN, GAE and GAT) are unable to efficiently build high-quality representations of regions with sparse crime data via the spatial-temporal graph.
IV-F Hyperparameter Study (RQ4)
@@
We conduct the hyperparameter study on the parameters including the mask ratio of nodes and links with the range $[0.1,0.3,0.5,0.7]$, and the number of CGN layer from the range $[2,3,4,5]$ on crime prediction task and traffic predition  w.r.t New York and Chicago in terms of MAE, MAPE and RMSE. We provide results in Figure 5. From figures, we can have the following observations: (1) For the impact of mask ratio of nodes and structure, we find that the model STGMAE can achieve the best performance on crime prediction and traffic prediction on all datasets when the mask ratio of nodes and links is set as 0.7. And the performance decreases with the decrease of mask ratio. This verifies that masking nodes and links do not always bring better model representation ability. Smaller mask ratios (e.g., $[0.1,0.3,0.5]$) may not denoise nosiy nodes and links completely, which leads to bad performance. (2) For the effect of the number of GCN layer, we find that our framework STGMAE achieves the best performance when the number of GCN layer is set as 2. We notice that the performance decreases with the increasing number of GCN layer. This may be caused by over-smoothing issue on region representation with the increasing number of the GCN layer. (3) We also investigate the consuming time of our model STGMAE with varied hyperparameters. We observe that time cost increases with increasing mask ratio of nodes and links, since our model STGMAE requires to reconstruct more nodes’ representation and new links when masking more nodes and links. Similarity, time consuming is larger when the more layers of GCN due to that the message passing between GCN layers need more time.
IV-G Model Efficiency Study (RQ5)

### v2

@@ 3.1.3. Implementation Details
3.1.4. Baselines
We use 15 baselines from three research lines namely, graph representation approaches, graph contrastive learning models, and spatial-temporal region representation methods for comparison in order to thoroughly assess our HGAurban model.
@@ 3.1.4. Baselines
  • Node2vec (Grover and Leskovec, 2016) encodes graph structural data using a random walk-based skip-gram.
  • GCN (Kipf and Welling, 2017) carries out the convolution-based message transfer along the edges between neighboring nodes for refinement.
  • GraphSage (Hamilton et al., 2017) is a graph neural architecture that permits the accumulation of information from the sub-graph structures.
  • GAE (Kipf and Welling, 2016) maps node into latent embedding space by the Graph Auto-encoder using the input reconstruction objective.
  • GAT (Veličković et al., 2018) varies the degrees of significance among nearby nodes, and improves GNNs’ capacity to discriminate between diverse inputs. To optimize the performance of GNN-based approaches, the number of information propagation layers is adjusted from a range of {1,2,3,4}.
@@
  • GraphCL (You et al., 2020) is based on the maximizing of mutual knowledge and generates many contrastive viewpoints for augmentation.
  • RGCL (Li et al., 2022) have been fine-tuned to manage the contrastive augmentation in order to allow for fair comparison.
@@
  • POI (Rahmani et al., 2019) uses TF-IDF tokenization with the provided POI matrix and the POI properties to represent spatial regions.
  • HDGE (Wang and Li, 2017) creates a crowd flow graph using human trajectory data and embeds regions into latent vectors to protect the structural data of the network and capture mobility patterns.
  • ZE-Mob (Yao et al., 2018) uses region correlations to create embeddings while taking into account human movement and taxi moving traces based on multi-view data sources.
  • MV-PN (Fu et al., 2019) models intra-regional and inter-regional correlations with an encoder-decoder network.
  • CGAL (Zhang et al., 2019) is a technique for graph-regularized adversarial learning that uses pairwise graph relationships to embed regions in latent space via the graph attention neural network.
  • MVURE (Zhang et al., 2021) makes use of the graph attention mechanism to simulate region correlations with the built-in properties of the region and data on human mobility.
  • AutoST (Zhang et al., 2023a) is a method that utilizes contrastive self-supervised learning method via multi-view graph to model region representation learning problem on traffic and crime prediction tasks.
  • MGFN (Wu et al., 2022) aggregates information for both intra-pattern and inter-pattern patterns, and encodes region embeddings with multi-level cross-attention without region function information.
@@
Figure 2. Results on NYC and CHI crime for four crime types w.r.t different data density degrees.
3.2. Effectiveness Evaluation (RQ1)
Table 2 shows a performance comparison of all methods on three tasks (crime prediction, traffic prediction, and house price prediction) in terms of MAE, MAPE, and RMSE.
Crime prediction. We show crime prediction results in Table 2. From Table 2, we summarize observations as follows:
  • On both datasets, our framework HGAurban performs best at predicting crimes, indicating the viability of our generative spatio-temporal graph self-supervised learning approach. In specific, we attribute the benefits to: i) HGAurban can capture both the mutli-view area relationships from multi-typed data sources (i.e., POI semantics, human motion traces, geographical coordinates) thanks to masking different types of relations. ii) The adaptive data augmentation way of HGAurban via the mask mechanism on nodes and relations improves the representation ability against noise disturbance and skewed data distribution.
  • While existing approaches for region representation (such as CGAL, MVURE, and MGFN) aim to capture the mobility-based region correlations, they are susceptible to noisy edges in their built mobility-based region graphs. The performance of region representation learning will be hampered by the propagation of information propagation between less-relevant regions. Besides, the spatial dependence modeling in such baselines is constrained by the skewed distribution of mobility data. Our HGAurban   which is complementary to the mobility-based region connections, adds region self-discrimination supervision signals to the region embedding paradigm to solve these limitations.
  • The large enhancement of HGAurban over the compared graph embedding techniques (i.e., GCN and GAT) further supports the significance of generative self-supervised learning paradigm. When a message passes over the edges of a task-irrelevant region in the graph neural paradigm, noise effects will be enhanced. The information compiled from all the nearby mobility-based or POI-based graph regions may be misleading when it comes to encoding the actual underlying crime trends.
Traffic prediction. We present traffic prediction results in Table 2 with the following observations:
  • The significance of the generative self-supervised learning with the mask mechanism is further supported by the significant performance improvement of HGAurban over the compared graph embedding techniques (such as GCN, GAT, and GAE). In the graph neural paradigm, noise effects will be amplified when messages crosses edges of regions that are unrelated to the task.
@@ 3.2. Effectiveness Evaluation (RQ1)
Figure 3. Hyperparameter study on traffic prediction
Refer to caption Refer to caption
Figure 4. Ablation study of HGAurban on traffic prediction
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 5. Ablation study of HGAurban on crime prediction
3.3. Ablation Study (RQ2)
@@
To evaluate the effectiveness of each component of our framework HGAurban, we perform an ablation study and present results w.r.t traffic prediction and crime prediction in Figure 4 and Figure 5. Particularly, we conduct an ablation study several variants including “RP GCN” by replacing GCN with MLP, “w/o Structure Mask”, “w/o Node Mask”, and “w/o Edge Type Mask” by removing masking links or masking nodes.
@@
Analysis on Crime prediction. The figure demonstrates the impact of different components on region representation for crime prediction. “RP GCN” replacing GCN with a 2-layer MLP, contributes the most to performance improvement compared to “w/o Structure Mask”, “w/o Edge Type Mask”, and “w/o Node Mask” GCN captures spatial-temporal graph structure semantics, unlike MLP. “w/o Node Mask” has a larger impact than “w/o Structure Mask” as masking nodes directly affects region representations. Masking different link types positively affects region representations, emphasizing the significance of masking links in the heterogeneous graph. “w/o Edge Type Mask” which doesn’t mask edges, shows a notable performance decline, highlighting the effectiveness of edge type masking in our HGAurban model. Overall, integrating GCN, node masking, and edge-type masking enhances the performance of our HGAurban framework.
@@
Analysis on Traffic prediction. Based on Figure 4, GCN has the greatest impact on improving region representations and traffic prediction, especially for longer time horizons. Masking edge types significantly improves traffic flow prediction accuracy. Masking nodes and graph structure also contribute positively by refining region representations and capturing spatial relationships. Incorporating POI information enhances region representations and enables the encoding of different region relations. Overall, the combination of GCN, edge type masking, node masking, and POI integration contributes to the success of our HGAurban method in learning region representations and achieving accurate traffic predictions.
3.4. More Effectiveness Evaluation Results (RQ1)
@@
Results for the data on home prices in Chicago and New York City are shown in Table 2 and more analysis details are below. A Lasso regression approach uses the input embeddings from the learned region representations created using various methods.
@@
  • HGAurban consistently performs at the highest level. On the other hand, MVURE and MGFN perform better when using data from the housing markets of New York and Chicago, respectively. This is because they mine human patterns, which connect comparable or nearby places. We also notice that the POI technique performs well because it creates connections between POIs of regions and core shopping regions, which include POIs that are distinct from those in remote towns.
@@
  • Table 3 shows results of improvement of existing widely used traffic prediction methods (e.g., GMAN, STGCN) based on several region representation baselines and our method HGAurban. The superior performance further validates the effectiveness of generative self-supervised learning with a masking mechanism for denoising data. In contrast, methods like MVURE and MGFN require significantly more data to uncover human patterns.
Table 3. Overall Performance Comparison in Traffic Prediction Improvement
@@
w/HGAurban 1.8403 2.7032 w/HGAurban 1.2415 3.6375
3.5. Model Robustness Study (RQ3)
@@
We also conduct experiments to test how well our framework HGAurban stands up to data sparsity. To accomplish this, we assess the forecast accuracy of regions with various densities independently. Here, the ratio of non-zero elements (crime occurs) in the region-specific crime occurrence sequence $\mathbf{Z}_{r}$ is used to estimate the density degree of each region. We divide sparse areas with a crime density degree of less than 0.5 into two groups, (0.0, 0.25] and (0.25, 0.5] specifically. Figure 2 presents the evaluation findings. We see that our HGAurban constantly performs better than alternative approaches. Thus, this experiment once more indicates how masking mechanism on nodes and the structure for self-supervised learning into the spatial-temporal region representation learning provides very accurate and reliable crime prediction in all scenarios with varying crime density degrees. To mine human behaviors or movement patterns, existing region representation algorithms (such MVURE and MGFN) require a huge quantity of data, which results in lower performance on the sparse data. Another finding acknowledges that typical graph learning techniques now in use (such as GNN-based approaches including GCN) are unable to efficiently build high-quality representations of regions with sparse crime data via the spatial-temporal graph.
@@
The sencond line of Figure 2 present more details about the evaluation findings of the model robustness. We see that HGAurban constantly performs better than alternative approaches. Thus, this experiment once more indicates how masking mechanism on nodes and the structure for self-supervised learning into the spatial-temporal region representation learning provides very accurate and reliable crime prediction in all scenarios with varying crime density degrees. To mine human behaviors or movement patterns, existing region representation algorithms (MVURE and MGFN) require a huge quantity of data, which results in lower performance on the sparse data. Another finding acknowledges that typical graph learning techniques now in use (i.e., GNN-based approaches like CGN) are unable to efficiently build high-quality representations of regions with sparse crime data via the spatial-temporal graph.
3.6. Hyperparameter Study (RQ4)
@@
We conduct the hyperparameter study on the parameters including the mask ratio of nodes and links with the range $[0.1,0.3,0.5,0.7]$, and the number of CGN layer from the range $[2,3,4,5]$ on crime prediction task and traffic predition  w.r.t New York and Chicago in terms of MAE, MAPE and RMSE. We provide results in Figure 3. From figures, we can have the following observations: (1) For the impact of mask ratio of nodes and structure, we find that the model HGAurban can achieve the best performance on crime prediction and traffic prediction on all datasets when the mask ratio of nodes and links is set as 0.7. And the performance decreases with the decrease of mask ratio. This verifies that masking nodes and links do not always bring better model representation ability. Smaller mask ratios (e.g., $[0.1,0.3,0.5]$) may not denoise nosiy nodes and links completely, which leads to bad performance. (2) For the effect of the number of GCN layer, we find that our framework HGAurban achieves the best performance when the number of GCN layer is set as 2. We notice that the performance decreases with the increasing number of GCN layer. This may be caused by over-smoothing issue on region representation with the increasing number of the GCN layer. (3) We also investigate the consuming time of our model HGAurban with varied hyperparameters. We observe that time cost increases with increasing mask ratio of nodes and links, since our model HGAurban requires to reconstruct more nodes’ representation and new links when masking more nodes and links. Similarity, time consuming is larger when the more layers of GCN due to that the message passing between GCN layers need more time.
3.7. Model Efficiency Study (RQ5)